### v2

# STELAR-VISION: Self-Topology-Aware Efficient Learning for Aligned Reasoning in Vision

Abstract
Vision-language models (VLMs) have made significant strides in reasoning, yet they often struggle with complex multimodal tasks and tend to generate overly verbose outputs. A key limitation is their reliance on chain-of-thought (CoT) reasoning, despite many tasks benefiting from alternative topologies like trees or graphs. To address this, we introduce STELAR-Vision, a training framework for topology-aware reasoning. At its core is TopoAug, a synthetic data pipeline that enriches training with diverse topological structures. Using supervised fine-tuning and reinforcement learning, we post-train Qwen2VL models with both accuracy and efficiency in mind. Additionally, we propose Frugal Learning, which reduces output length with minimal accuracy loss. On MATH-V and VLM_S2H, STELAR-Vision improves accuracy by 9.7% over its base model and surpasses the larger Qwen2VL-72B-Instruct by 7.3%. On five out-of-distribution benchmarks, it outperforms Phi-4-Multimodal-Instruct by up to 28.4% and LLaMA-3.2-11B-Vision-Instruct by up to 13.2%, demonstrating strong generalization. Compared to Chain-Only training, our approach achieves 4.3% higher overall accuracy on in-distribution datasets and consistently outperforms across all OOD benchmarks.
Datasetshttps://huggingface.co/collections/Stellar-Neuron/stelar-vision-collection
Project Pagehttps://stellar-neuron.github.io/stelar-vision/
1 Introduction
Recent advances in large language models (LLMs) have significantly improved reasoning capabilities, with models like GPT-o3 achieving strong performance on complex mathematical and scientific tasks. This progress has extended into the multimodal domain through vision-language models (VLMs) such as GPT-4o (Hurst et al. 2024), GPT-4o-mini (OpenAI 2024), and Qwen2.5-VL (Bai et al. 2025). Despite the recent advances, there is still room of improvement in open-sourced VLMs when tackling complex vision-based reasoning tasks (e.g., math and science questions), and the path to enhance their abilities under an affordable training budget remains under-explored.
To address this, we begin by analyzing VLMs’ reasoning behaviors and find that the popular models, both open-source and closed-source, tend to default to the chain-of-thought (CoT) (Wei et al. 2023) generation. This behavior reflects the prevailing trend in their training data, which is overwhelmingly dominated by CoT-style reasoning samples. However, our empirical analysis reveals that different questions benefit from different reasoning topologies, such as Chain, Tree, or Graph structures (Figure 1). The benefits of diverse reasoning topologies have yet been well studied or effectively incorporated into existing training pipelines. Moreover, even state-of-the-art reasoning VLMs tend to generate unnecessarily verbose responses, i.e., “overthinking”, which increases the computational cost and makes real-time applications less viable. We find that there is a correlation between the topological reasoning structures and the output sequence length, thus providing an insight to the overthinking problem created by the CoT reasoning.
We introduce Self-Topology-Aware-Efficient-Learning for Aligned Reasoning in Vision, STELAR-Vision, a training framework for topology-aware vision-language reasoning. The central to this framework is TopoAug, a synthetic data generation pipeline that produces augmented reasoning topologies with diverse structures, including $\{Chain,Tree,Graph\}$. Specifically, we generate a set of question-answer responses, where each question is repeatedly answered by different topological reasoning structures. For each question, we assign a preferred topology through an automated annotation process. The generated responses and their labels are used to post-train VLMs via supervised fine-tuning (SFT) and Reinforcement Learning (RL) (Meng, Xia, and Chen 2024). Regarding the learning process, we observe that RL further amplifies the performance gains introduced by topological augmentation. Moreover, we propose to further increase efficiency with Frugal Learning by promoting shorter responses. We show that only by leveraging the benefits of augmented reasoning topologies, can the model generate concise responses and maintain high accuracy with minimal performance drops, while the counterpart model trained only on CoT-style data fails to gain efficiency without incurring a greater loss in accuracy.
Figure 1: Limitations of the Popular Chain-of-Thought Reasoning Structures. The widely adopted Chain-of-Thought (CoT) reasoning paradigm (in green) often results in unnecessarily verbose reasoning processes, as demonstrated in the first example. Under CoT reasoning, the model redundantly counts each cube, whereas with GraphGraph topology (in blue), it quickly identifies the key point of the question. In the bottom-row example, CoT reasoning begins with a detailed examination of each subplot but ultimately arrives at an incorrect answer. In contrast, TreeTree topology (in red) initiates reasoning with a high-level overview before delving into specific features. In both scenarios, CoT-style reasoning proves suboptimal.
The augmented topologies expand the exploration space, allowing RL to discover higher-quality optima and leading to superior performance improvement. In contrast, the model trained with CoT-style data only sees diminishing returns. We also test our models on five out-of-distribution datasets, and it consistently outperform its base model, while consuming fewer tokens. which suggests strong generalization of our proposed STELAR-Vision.
Our contributions are summarized as follows:
  • We propose STELAR-Vision, a training framework explicitly trained for topology-aware reasoning. It leverages diverse reasoning topologies such as chains, trees, and graphs, aligns reasoning paths with question characteristics, and enables adaptive and efficient multimodal inference.
  • We introduce TopoAug, a data generation pipeline that automatically produces diverse topological reasoning and annotates optimal structures per question. We also integrate Frugal Learning into the learning framework, achieving reductions in output length with minimal accuracy tradeoff.
  • By conducting experiments with post-training supervision and reinforcement learning, STELAR-Vision improves accuracy by 9.7% over its base model and its larger variant Qwen2VL-72B-Instruct by 7.3%. On the out-of-distribution dataset, it surpasses The Frugal Learning variant reduces output length by 18.1% while maintaining comparable accuracy.
2 Related Work
2.1 Topological Reasoning in Language and Vision Models
Chain-of-Thought (CoT) prompting (Wei et al. 2023) is a widely used reasoning strategy in LLMs and VLMs, guiding models to generate step-by-step solutions. However, its linear structure may not suit all tasks. To address this, Tree-of-Thought (ToT) (Yao et al. 2023) enables branching exploration, while Graph-of-Thought (GoT) (Besta et al. 2024) supports iterative and global reasoning. Both improve performance on complex tasks like TSP, algorithmic problem-solving, and multi-stage decision-making.
These methods, however, often rely on rule-based topology generations through prompt engineering or sampling, and are limited to language-only settings. In contrast, our framework automatically generates diverse topological structures and trains a VLM to adaptively select the optimal one per instance during decoding, enabling more flexible and generalizable reasoning.
2.2 Reinforcement Learning for LLM and VLM Reasoning
Reinforcement learning (RL) is a key technique for aligning LLMs and VLMs with desired behaviors in reasoning and preference modeling. Approaches like RLHF (Stiennon et al. 2022; Ouyang et al. 2022) and Constitutional AI (Bai et al. 2022) enable models to acquire complex reasoning and ethical behaviors via preference optimization.
Reward-based methods such as PPO (Schulman et al. 2017), RPO (Yin et al. 2024), and GRPO (Shao et al. 2024) rely on explicit rewards, while reward-free approaches like DPO (Rafailov et al. 2024), SimPO (Meng, Xia, and Chen 2024), and ORPO (Hong, Lee, and Thorne 2024) achieve comparable results without reward modeling. These methods are widely applied to mathematical reasoning, long-horizon tasks, and instruction tuning.
In VLMs, RL has been used to enhance structured reasoning and safety-critical applications. VLM-RL (Huang et al. 2024) improves decision-making in autonomous driving, MedVLM-R1 (Pan et al. 2025) ensures safety in medical imaging, and RLVR (Chen et al. 2025) boosts OOD generalization in tasks like visual counting and open-ended QA.
Building on these insights, we show that combining topology-aware data generation with RL (e.g., SimPO) improves both accuracy and efficiency. Topological diversity expands the exploration space, increasing the likelihood of discovering stronger reasoning strategies during RL.
2.3 Curriculum Learning and Structured Reasoning
Curriculum learning (Bengio et al. 2009) trains models on progressively harder tasks, aiding structured reasoning. Recent works extend this idea: Xi et al. (Xi et al. 2024) use reverse curricula for LLMs, Zhao et al. (Zhao et al. 2024) propose Auto-CEI for ability-aligned sampling, and Ma et al. (Ma, Jiang, and Huang 2025) design logic-guided curricula for in-context learning. Fine-tune-CoT (Ho, Schmid, and Yun 2023) learns CoT from larger reasoning models.
In VLMs, LlamaV o1 (Thawakar et al. 2025) schedules visual examples for step-wise reasoning, and VLM-R1 (Shen et al. 2025) integrates curriculum-aware GRPO for stability and interpretability. However, most curricula are static and predefined.
In contrast, our approach is an end-to-end learning via a dataset from TopoAug, where synthetic reasoning paths evolve with model performance. Coupled with RL and topological diversity, our training framework promotes structured reasoning in a more scalable and generalizable way.
3 Method
In this section, we first construct topology-aware responses on two mathematical datasets. We then investigate the relationship between the topological reasoning and response accuracy. Finally, we present the topology-aware training framework shown in Figure 2.
Figure 2: An overview of the STELAR-Vision framework
3.1 Constructing Topology-Aware Responses
Data
We use two math datasets: MATH-V (Wang et al. 2024a) (3,040 visual problems) and VLM_S2H (Park et al. 2025) (7,000 logic puzzles), each pairing images with questions, giving diverse samples requiring high-level reasoning
TopoAug
We generate Topology-Aware responses using topologies $T=\{Chain,Tree,Graph\}$, prompted via Qwen2-VL-7B-Instruct (Wang et al. 2024b) and GPT-4o-Mini (OpenAI 2024a) with extensive degrees of freedom in maximum depth, number of children, and number of neighbors. Please see the supplementary for detailed prompts.
Topology and Outcome Labels
Each response $r$ has an Outcome Label $\mathcal{H}_{r}\in\{0,1\}$, and is assigned label 1 if correct and 0 otherwise. Each question-topology pair gets a Topology Label $\mathcal{F}_{q,t}=\frac{N_{\text{correct}}(q,t)}{N_{\text{total}}(q,t)}$ based on accuracy, where $N_{\text{correct}}(q,t)$ is the number of correct responses using topology $t$ for question $q$, and $N_{\text{total}}(q,t)$ is the total number of responses generated using $t$.
Problem Difficulty Segmentation
Problems are labeled Easy (>>85th percentile), Hard (<<15th), or Medium based on topology score distributions.
3.2 Analysis: Topological Reasoning Structures
We first show that different questions are better solved using distinct reasoning topologies. We evaluate the two aforementioned vision-language models (VLMs). We exclude the newer Qwen2.5-VL(Qwen et al. 2025) series from our analysis due to its relatively unstable performance in generating diverse reasoning topologies.
We first prompt the models to generate responses using their default reasoning behavior. Next, we explicitly instruct them to reason using three distinct topologies: Chain, Tree, and Graph. We compute a topology-wise Win Rate to assess the performance of each topological reasoning structure, which is defined below. We then conduct a subject category-wise study on win rate.
Win Rate: We measure across the entire dataset and calculate the percentage of occurrence where a topology $t$ is the best performing reasoning structure among the three topology types in Equation 1:
$\text{Win Rate}(t)=\frac{\sum_{q\in Q}\mathds{1}_{t}(\operatorname*{arg\,max}_{t^{\prime}\in T}\mathcal{F}_{q,t^{\prime}})}{N_{Q}}$ (1)
where $N_{Q}$ is the total number of questions, and $\mathds{1}$ is the indicator function. For each question, the topology with the highest Topology Label $\mathcal{F}_{q,t}$ wins. We first analyze which topological reasoning structure works the best for the questions. The Win Rate statistics is presented in Table 1.
Chain Tree Graph
Win Rate 49% 28% 23%
Table 1: Comparison of Win Rates across reasoning topologies.
Figure 3: Comparison of topology accuracy across subjects: Accuracy of ChainChain, TreeTree, and GraphGraph reasoning topological structures per subject of MATH-V dataset. ChainChain remains the best overall reasoning structure, while TreeTree, and GraphGraph perform better in at reasoning subjects such as “graph theory” or “statistics”.
Our analysis in Table 1 reveals that $Chain$ structure reasoning remains the best overall reasoning pattern in both VLMs, regardless of model size. However, $Tree$ and $Graph$ collectively account for more than half of the winning responses. We conduct a detailed investigation of each topology’s performance across subject domains by following the 10 subject categories in the MATH-V dataset and measuring the accuracy of each reasoning topology per subject, as illustrated in Figure 3. And thus we arrive at a conclusion that different problems benefit from different topologies, and identifying the optimal ones yields accuracy gains. This aligns with theoretical expectations—while straightforward tasks are adequately addressed by CoT-style reasoning, more complex or structurally intricate problems require richer topological representations to capture their underlying relationships effectively.
Topology-Wise Generation Length We further investigate the reasoning generation token length distribution within the TopoAug dataset. As illustrated in Figure 4, while most generations exhibit an average length of approximately 550 tokens, the $Chain$ topology produces the longest generations with a right-skewed distribution that favors extended reasoning processes. In contrast, both $Tree$ and $Graph$ topologies demonstrate similar token length distributions with overall a lower length, indicating comparable reasoning verbosity across these structured approaches.
Figure 4: Distribution of generated reasoning token length of ChainChain, TreeTree, and $Graph$ topological structures in TopoAug Dataset. The box within each violin plot represents the median, and 25% and 75% percentile thresholds.
3.3 STELAR-Vision Post-Training
Our finding propels us to make two reasonable assumptions:
  • Assumption 1: A model trained with topologically diverse reasoning structures—without increasing data volume—can achieve higher reasoning accuracy by learning to adaptively identify the best reasoning structure for each problem at test time.
  • Assumption 2: Building up Assumption 1, a learning mechanism can be designed to encourage concise yet accurate outputs, enhancing inference efficiency with minimal performance loss.
We device a Post-training framework that consists of two phases: Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL).
Phases 1: Supervised Fine-tuning
We collect the data generated via TopoAug and combine it with three additional datasets, OKVQA(Marino et al. 2019), A-OKVQA(Schwenk et al. 2022), and LLaVA150k-Instruct (Liu et al. 2023). We include them as general VQA tasks to preserve the model’s basic VQA capabilities. Note that the additional 3 VQA datasets are used without any topological augmentation to avoid altering their original structure and maintain generalization. We perform SFT on TopoAug-generated data mixed with general VQA datasets, using a three-step filtering process: (1) balanced sampling from Easy, Medium, and Hard problems, (2) keeping only responses with positive outcome labels, and (3) rejection sampling with a 7B Outcome Reward Model (ORM) (Ouyang et al. 2022), trained with both topology and outcome labels to select higher-quality samples with greater win potential.
The fine-tuning uses LoRA (Hu et al. 2021) with next-token prediction (NTP), minimizing the loss
$\mathcal{L}_{\text{NTP}}=-\sum_{t=1}^{T}\log P_{\theta}(y_{t}\mid y_{<t},x)$ (2)
where $x$ is the input, $y=(y_{1},\dots,y_{T})$ the target, and $P_{\theta}$ the model’s predicted token probabilities.
Phases 2: Reinforcement Learning
We follow prior work (Chu et al. 2025; Zhai et al. 2024) by initializing reinforcement learning (RL) from an SFT checkpoint, which improves both in-distribution accuracy and OOD generalization. We adopt SimPO (Meng, Xia, and Chen 2024) for its simplicity and alignment with test-time behavior. Its objective encourages preferred responses over less-preferred ones, and is defined as
$\begin{split}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=&-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Bigg[\log\sigma\Bigg(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}\mid x)\\ &\qquad\qquad\qquad\quad-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}\mid x)-\gamma\Bigg)\Bigg]\end{split}$ (3)
where $x$ is the input, $y_{w}$ and $y_{l}$ are the preferred and less-preferred responses, $\pi_{\theta}$ is the policy, $\sigma$ is the sigmoid, and $\beta$, $\gamma$ are temperature and margin parameters.
We compare two RL setups: one trained on TopoAug-based preference pairs and one on Chain-only data of equal size. Correct responses are treated as the preferred responses. To prevent data leakage, we remove topology prompts during training, so the model must infer optimal structures at test time.
Model In-Distribution Accuracy (%) Out-of-Distribution Accuracy (%)
VLM_S2H MATH-V Overall Geometry3K We-Math PolyMath SciBench LogicVista
GPT-4o (OpenAI 2024b) 32.0 28.0 30.7 57.0 66.4 25.0 31.1 34.6
LLaVA-v1.6-Mistral-7B (Liu et al. 2024) 26.0 8.0 18.0 20.6 26.0 9.2 3.4 18.5
Llama-3.2-11B-Vision-Instruct (Meta 2024) 22.0 10.0 18.0 35.0 37.8 22.2 10.7 24.8
MiniCPMv2.6-8B (Yao et al. 2024) 1.5 13.0 18.7 45.0 50.2 14.4 8.5 20.7
Phi-4-multimodal-5.6B-instruct (Abouelenin et al. 2025) 23.0 11.0 22.0 8.4 35.8 10.2 10.2 6.7
InternVL3-9B (Zhu et al. 2025) 25.0 21.0 27.3 41.2 51.4 21.6 20.3 32.6
Qwen2VL-72B-Instruct (Yang et al. 2024) 21.0 20.0 20.7 50.2 60.6 13.0 25.4 28.8
Qwen2VL-7B-Instruct (Yang et al. 2024) 21.0 13.0 18.3 35.2 46.6 16.0 10.7 17.0
Chain-Only 25.0 21.0 23.7 31.4 42.2 17.2 10.7 25.4
STELAR-Vision-SFT 28.0 24.0 26.7 44.4 47.4 24.8 9.0 33.3
STELAR-Vision-RL-ONLY 24.0 23.0 23.7 32.8 39.0 26.0 17.5 23.9
STELAR-Vision 31.0 22.0 28.0 36.8 51.0 23.8 12.4 29.0
Table 2: Quantitative Evaluation. STELAR-Vision achieves strong gains across both in-distribution and out-of-distribution reasoning benchmarks. On ID datasets, it outperforms its base model Qwen2VL-7B-Instruct by 9.7%, and even surpasses the larger Qwen2VL-72B-Instruct by 7.3%. On OOD benchmarks, it exceeds Phi-4-multimodal-instruct by up to 36% and LLaMA-3.2-11B-Vision-Instruct by up to 13.2%. Compared to Chain-Only training, STELAR-Vision achieves up to 13% higher accuracy, highlighting the power of topological augmentation.
Frugal Learning
While recent work has explored efficient reasoning in LLMs (Arora and Zanette 2025; Aggarwal and Welleck 2025), these approaches rely on additional reward models, lack cost-controllability, and do not focus on vision-based reasoning. To improve test-time reasoning efficiency, we propose Frugal Learning, which trains a compact variant called STELAR-V-Short. We propose and compare two training strategies:
Variant 1: STELAR-Vision-Short. We first introduce a filter that identifies “short and correct” responses as preferred targets, where the outcome Label $\mathcal{H}_{r}=1$ and the generated token length falls below a 25% percentile threshold. The model undergoes training with both supervised fine-tuning (SFT) and reinforcement learning (RL). During RL training, we designate “short and correct” responses as winners and incorrect responses as losers in the preference optimization process. We also train a model with Chain-Only data with the same process and we denote it as Chain-Only-Short.
Variant 2: STELAR-Vision-Short. A natural alternative approach to encourage shorter and more efficient responses involves explicitly penalizing lengthy outputs. Building upon STELAR-Vision-Short, we investigate an alternative configuration, STELAR-Vision-Short, which maintains the preference for “short and correct” responses while treating both “incorrect responses” and “correct yet lengthy” responses as losers during RL preference training. This design encourages the model to avoid excessive token generation while maintaining accuracy in reaching the correct final answer.
4 Experiments
4.1 Experimental Setup
Datasets
We use In-Distribution (ID) datasets from the Method Section 3, splitting them into train/test set. We include 85K training samples from MATH-V and 160K samples from VLM_S2H. OKVQA and A-OKVQA provide 18k and 20K VQA pairs for training, with A-OKVQA adding rationales to training. We also use 17k instruction-following examples from LLaVA-150k for multimodal training. performed on the test sets of MATH-V and VLM_S2H.
For Out-of-Distribution (OOD) evaluation, we use five recent benchmarks with three math oriented datasets, Geometry3K (Lu et al. 2021), We-Math (Qiao et al. 2024) and PolyMath (Gupta et al. 2024), a STEM dataset SciBench (Wang et al. 2024c) and a generic logic reasoning dataset LogicVista (Xiao et al. 2024). In Table 3, we show the evaluation datasets as well as their respective sizes. Collectively, we gather 2,425 evaluation samples.
Models
Our base model is Qwen2VL-7B-Instruct (Wang et al. 2024b), chosen for its stable generation of all three topologies in 74% of cases. We exclude the newer Qwen2.5VL-7B-Instruct (Team 2024) for its instability in producing $Tree$ and $Graph$ structure reasoning. We also compare our models with additional open-source and proprietary models.
Dataset Subject Question Type Sample Size
VLM_S2H Math multiple-choice 200
MATH-V Math free-form, multiple-choice 100
Geometry3K Math multiple-choice 500
We-Math Math multiple-choice 500
PolyMath Math multiple-choice 500
SciBench STEM free-form 177
LogicVista Generic multiple-choice 448
Total 2425
Table 3: Summary of Evaluation Datasets. VLM_S2H and MATH-V are in-distribution datasets, while others are out-of-distribution
Evaluation Metrics
We report Accuracy as the main metric: $\text{Accuracy}=\frac{\sum_{q\in Q}\text{Answer}(q)=\text{GT}(q)}{N_{Q}}$, where GT stands for the ground-truth answer. To assess efficiency under Frugal Learning, we also consider Generated Token Length as a metric for reasoning efficiency.
All experiments are conducted on eight NVIDIA A100/H100 GPUs with 80GB of memory each. Supervised fine-tuning (SFT) on a 7B model with  50K–60K samples takes approximately 5–7 hours, while reinforcement learning (RL) on the same scale requires 8–10 hours, depending on system variability.
4.2 Overall Evaluation Results
We use STELAR-Vision to denote models trained with both phases of post-training, the -SFT suffix indicates models trained with supervised fine-tuning only, and -RL-ONLY indicates reinforcement learning directly from the base model without SFT. Results are averaged across 3 random seeds.
Table 2 shows results on the in-distribution datasets MATH-V and VLM_S2H as well as OOD datasets. STELAR-Vision achieves the highest in-distribution accuracy, significantly outperforming its base model Qwen2VL-7B-Instruct by 9.7%, and surpassing larger model LLaMA-3.2-11B by 10.0% and Qwen2VL-72B-Instruct by 7.3%.
Results on out-of-distribution (OOD) benchmarks highlight the strong generalization of STELAR-Vision. It outperforms its base model Qwen2VL-7B-Instruct on all five OOD datasets and trails GPT-4o by only 1.2% on PolyMath. Compared to recently popular model Phi-4-Multimodal-instruct, STELAR-Vision achieves significantly higher accuracy across all OOD tasks, including a +28.4% gain on the spatial reasoning benchmark Geometry3K. It also surpasses other strong open-source models such as LLaVA-v1.6, LLaMA-3.2-11B-Vision, MiniCPMv2.6-8B, and InternVL3-9B on most benchmarks, demonstrating the impact of our TopoAug-enhanced training. STELAR-Vision also outperforms baseline on LogicVista, which includes non-math tasks like 26.4% diagram and 18% spatial.
Interestingly, the -SFT and -RL-ONLY variants sometimes outperform the full model. We attribute this to: (1) -SFT potentially overfitting due to memorization (Chu et al. 2025), and (2) -RL-ONLY occasionally lacking alignment. Topological augmentation appears to mitigate both limitations by improving generalization and alignment.
Finally, we find that Chain-Only models, despite moderate in-distribution gains, fall short on OOD tasks, further validating the necessity of topology-aware training. See Section 4.3 for details.
4.3 Ablation Studies
We perform ablation studies comparing our models to counterparts trained solely on chain-based reasoning data. As shown in Table 4, STELAR-Vision consistently outperforms Chain-Only variants on in-distribution (ID) datasets, improving accuracy from 23.7% to 28% (+4.3%). Table 2 further confirms this trend across benchmarks. On out-of-distribution (OOD) datasets, STELAR-Vision achieves up to 8.8% higher accuracy than Chain-Only models.
These findings collectively demonstrate that STELAR-Vision benefits not only from data distillation or pattern memorization but also from a genuine ability to adaptively select optimal reasoning topologies based on the task.
Model VLM_S2H MATH-V Overall
Qwen2VL-7B-Instruct 21.0 13.0 18.3
Chain-Only-SFT 18.5 19.0 18.7
Chain-Only-RL-ONLY 23.5 15.0 20.7
Chain-Only 25.0 21.0 23.7
STELAR-Vision-SFT 28.0 24.0 26.7
STELAR-Vision-RL-ONLY 24.0 23.0 23.7
STELAR-Vision 31.0 22.0 28.0
Table 4: Impact of TopoAug Dataset and Training Methods. We present an ablation study on the in-distribution VLM_S2H and Math-V datasets to compare the performance of our models against counterparts trained exclusively on chain-based reasoning data across all training methods. STELAR-Vision consistently outperforms all Chain-Only variants across all ID datasets-specifically it improves the highest variant Chain-Only from 25% to 31% by 6%, and boosts overall accuracy by 4.3%, highlighting the effectiveness of topological augmentation.
Model Accuracy (%) Gen. Token Length
ID OOD
Qwen2VL-7B-Instruct 26.2 613.5 543.3
Chain-Only 28.7 878.4 742.6
Chain-Only-Short 23.9 843.1 713.0
STELAR-Vision-SFT 26.7 604.4 483.3
STELAR-Vision 31.6 556.7 523.4
STELAR-Vision-Short 28.7 455.7 498.6
STELAR-Vision-Short 21.9 538.9 555.9
Table 5: Comparison of accuracy and generated token length across models: STELAR-Vision improves performance while using fewer generation tokens. Frugal learning further improves generation efficiency.
4.4 Efficiency Gains from Frugal Learning
We assess the impact of Frugal Learning on both in-distribution (ID) and out-of-distribution (OOD) performance, as shown in Table 5. The difference between STELAR-Vision-Short and STELAR-Vision-Short is detailed in Section 3.3.
As shown in Table 5, STELAR-Vision-Short reduces token length by 101 (ID) and 24.5 (OOD), while still outperforming Qwen2VL-7B-Instruct by 2.5%. However, it shows a 2.9% accuracy drop compared to full STELAR-Vision. In contrast, STELAR-Vision-Short yields inconsistent length reduction and greater performance loss, likely due to conflicting optimization signals that penalize correct outputs.
Chain-Only-Short also shortens outputs but suffers from even larger accuracy drop. Notably, Chain-Only models fine-tuned with RL often generate overly verbose responses, even with Frugal Learning—highlighting the advantage of TopoAug in promoting concise and effective reasoning.
Model Dataset Tree Graph Chain
w/o STELAR-Vision Overall - - 100.00
w/ STELAR-Vision ID 14.3 9.7 76.0
We-Math 63.0 7.4 29.6
Geometry3K 96.4 3.0 0.6
LogicVista 22.7 15.6 61.7
PolyMATH 54.0 14.8 31.2
SciBench 54.2 23.2 22.6
Table 6: Impact of Training on Test-time Topology Selection. Percentage of reasoning topologies autonomously selected by each model on our evaluation datasets, without explicit prompting. ID denotes the in-distribution test split.
4.5 Why Our Method Works?
Our method achieves substantial improvements on both in-distribution and out-of-distribution benchmarks. We hypothesize two key reasons for this success. First, for models that already possess the ability to generate diverse reasoning topologies but lack the capability to select the optimal topology for a given problem (e.g., Qwen2VL-7B-Instruct), our approach enables the model to adaptively choose the most effective structure. Second, for models that might not inherently support diverse topologies, our framework could instill this ability through guided supervision and reinforcement learning, making our methodology broadly applicable. However, due to computational constraints, we will leave systematic study to isolate this effect for future work.
Empirical results support these hypotheses: STELAR-Vision consistently outperforms the Chain-only baseline across both in-distribution and out-of-distribution tasks, as shown in Table 2. Moreover, Table 6 reveals a notable post-training increase in the generation of tree and graph structures, which correlates with accuracy gains—demonstrating the utility of topological augmentation.
On out-of-distribution benchmarks, we observe that datasets such as LogicVista favor simpler, more generic reasoning and hence exhibit a higher frequency of chain-based reasoning. In contrast, datasets requiring more complex reasoning, such as Geometry3K and SciBench, show an increased prevalence of tree and graph topologies. This distribution indicates that the model does not merely memorize patterns but has genuinely learned to select the most suitable topology based on the nature of the problem.
5 Conclusion and Discussion
In this work, we propose STELAR-Vision, a training framework that enables VLM’s topology-aware reasoning ability via generated responses. STELAR-Vision enhances vision-language reasoning by leveraging diverse topological structures, achieving a 9.7% accuracy improvement over its base model and outperforming its larger variant Qwen2VL-72B-Instruct by 7.3%. The Frugal Learning reduces output length by 18.1% while maintaining comparable accuracy, surpassing Chain-Only baselines in both efficiency and task effectiveness. STELAR-Vision demonstrates strong generalization capabilities across OOD tasks, as evidenced by five diverse datasets spanning multiple domains. Compared to Chain-Only training, our method achieves 4.3% higher overall accuracy on in-distribution datasets and delivers consistent performance improvements across all OOD evaluations. These results highlight the benefits of topological augmentation and the model’s ability to adapt to a variety of reasoning challenges through dynamic reasoning structures.
Despite these promising results, STELAR-Vision relies on predefined topology types, and the dynamic relationship between problem structure and optimal reasoning topology remains underexplored. Future work will focus on enabling scalable, end-to-end topology induction and extending applicability to more advanced multimodal reasoning tasks.